# DiffSSC: Semantic LiDAR Scan Completion using Denoising Diffusion Probabilistic Models

### v2

@@ III-C Denoiser Design and Regularization
Figure 3: Architecture of our MinkUNet Denoiser. The conditional input, detailed in the red area, is inserted between each layer of denoiser to guide the generation of point cloud.
@@
As shown in Fig. 3, the denoiser is based on the MinkUNet architecture [24, 25, 26, 27]. Given the feature $\mathcal{F}$ extracted from a layer of MinkUNet, we integrate the conditional input and step information between layers to obtain the fused feature $\mathcal{F}^{\prime}$. The raw semantic point cloud $\mathcal{X}$ is encoded as a conditional input $\mathcal{C}$ using the same MinkUNet encoder. To embed the most relevant conditional input into the feature space, a closest point algorithm is employed to effectively align the conditional input with the features. Simultaneously, the step $t$ is encoded as $\tau$ using sinusoidal positional encodings. After passing through an MLP individually, the conditional input and step information are concatenated to form the weight $\mathcal{W}$. To align the dimensions with the feature $\mathcal{F}$, $\mathcal{W}$ is processed through an MLP to produce $\mathcal{W}^{\prime}$. Finally, $\mathcal{W}^{\prime}$ and $\mathcal{F}$ are element-wise multiplied to form the refined feature $\mathcal{F}^{\prime}$, which is then passed to the next layer.
As proposed by Nunes et al. [23], the regularization loss consists of a local term focusing on each point and a global term addressing the statistical characteristics. We extend this design to the semantic domain by employing an $L_{2}$ loss between the added noise and the model’s predictions, along with losses for the mean $\mu_{\epsilon}$ and variance $\sigma_{\epsilon}$ to regulate the global statistical properties. Thus, the overall loss is formulated as follows:
$L=L_{2}+\lambda(L_{\text{mean}}+L_{\text{var}})=L_{2}+\lambda(\mu_{\epsilon}^{2}+(\sigma_{\epsilon}-1)^{2})$ (5)
where $L_{\text{diff}}$ is the regularization term focused on local features, commonly used in DDPM models, while $L_{\text{mean}}=\mu_{\epsilon}^{2}$ and $L_{\text{var}}=(\sigma_{\epsilon}-1)^{2}$ are designed to ensure the overall noise distribution aligns with a Gaussian distribution.
@@ III-D Refinement
Inspired by Lyu et al. [30], we design a refinement and upsampling scheme based on MinkUNet to further enhance the density of the diffusion model’s output. This module predicts $k$ bias $b_{k}\in\mathbb{R}^{3}$ for each point position in the completed scene, while the semantics are propagated to the biased points. The refinement module offers a marginal improvement in scene quality, but it functions more like interpolating points in the gaps, rather than learning to predict missing geometry and semantics. The main contribution is made by the diffusion model, as will be demonstrated in the ablation study.
IV Experiment
IV-A Experiment Setup
IV-A1 Datasets
We evaluate our approach using the SemanticKITTI [31] and SSCBench-KITTI360 [32] datasets. SemanticKITTI is a widely used autonomous driving dataset that provides point-wise annotations on raw LiDAR point clouds, extending the KITTI dataset to semantic study. Additionally, it builds the SSC benchmark by accumulating annotated scans within sequences. SSCBench-KITTI360 is another SSC benchmark derived from KITTI-360 [33], featuring LiDAR scans encoded the same as SemanticKITTI. This consistency allows SSC methods evaluated on SemanticKITTI to be seamlessly transferred to the KITTI-360 scenario. However, these SSC benchmarks only use the front half of the LiDAR scan (180° LiDAR field-of-view (FoV)) as input, which is not ideal for LiDAR-centered point cloud data. To address this, we additionally incorporate the rear half of the point cloud, facilitating the evaluation of SSC approaches on LiDAR-centered data. Our model is trained and validated purely on SemanticKITTI, using sequences 00-06 for training and sequences 09-10 for validation. We evaluate our model using the official validation sets of both datasets: sequence 08 of SemanticKITTI and sequence 07 of SSCBench-KITTI360.
IV-A2 Training and Inference
To train DiffSSC on the 360 LiDAR FoV, we generate the ground truth following the guidelines of SemanticKITTI. First, given the pose of each frame, we construct the global map by aggregating the semantic LiDAR sweeps within the sequence. Next, we extract the neighboring region around the key frame, specifically, a spherical area centered on the LiDAR with a radius of 60 meters. The model is trained on an NVIDIA A6000 GPU for 20 epochs. For the diffusion parameters, we employ a cosine schedule to modulate the intensity of noise at each step. Specifically, we set $\beta_{0}=3.5\times 10^{-5}$ and $\beta_{T}=0.007$, with the number of diffusion steps $T=1000$, and define $\beta_{1},\dots,\beta_{T-1}$ using the following equation.
@@ IV-A2 Training and Inference
We also set the ratio of global regularization to $\lambda=5.0$.
IV-A3 Baselines
@@
We compare our approach against LMSCNet [12], JS3C-Net [15], and LODE [34]. Both LMSCNet and JS3C-Net take the front half of the quantized LiDAR sweep as input and are evaluated in the SSC benchmark of SemanticKITTI. LODE primarily focuses on geometry completion using implicit representation. However, to demonstrate its flexibility, the authors also report results with extended semantic parsing. A common limitation among these baselines is that they are trained on point clouds within a 180 LiDAR field of view. To fairly compare with our method, we split the 360 LiDAR point cloud input into two halves and feed them separately into the baselines. The outputs from these two halves are then concatenated to obtain a 360 result. Additionally, while these baselines have only been tested on SemanticKITTI, we also ran them on SSCBench-KITTI360 as a supplementary experiment. Since the semantic labels and overall pipeline in SSCBench-KITTI360 are consistent with SemanticKITTI, the baselines can be seamlessly applied to this dataset.
IV-A4 Evaluation Metrics and Pipeline
Despite our task being set in a 360 LiDAR point cloud completion context, we aimed to retain the baseline settings as closely as possible to ensure a fair comparison. In the raw SSC setting of SemanticKITTI, the scene is limited to a cuboid region, represented in the LiDAR’s local coordinate system as: $V_{\text{kitti}}=\{(x,y,z)\mid x\in[0,51.2]\text{ m},y\in[-25.6,+25.6]\text{ m},z\in[-3.2,+3.2]\text{ m}\}$, which corresponds to a region associated only with the LiDAR’s $[-90^{\circ},+90^{\circ}]$ FoV. To cover the full $[-180^{\circ},+180^{\circ}]$ panoramic range while preserving spatial symmetry, we selected an evaluation region within the LiDAR’s local coordinate system defined as: $V_{\text{ours}}=\{(x,y,z)\mid x\in[-51.2,51.2]\text{ m},y\in[-25.6,+25.6]\text{ m},z\in[-3.2,+3.2]\text{ m}\}$.
We directly used the baselines’ official code and checkpoints to predict the front and rear parts of the scene, with each part of the LiDAR sweep input separately. Although the baselines were trained only on the front part of the scene, the statistical characteristics of LiDAR data in the front and rear regions are similar, suggesting that a model trained using only the front half of the data remains effective for the rear region as well. Although our ground truth generation covers a spherical region with a radius of 60 meters centered on the LiDAR, we limited our evaluation to the region predicted by the baselines. Additionally, the unknown areas defined by the raw dataset were mapped into $V_{\text{ours}}$ using known poses, and these unknown areas were excluded from the evaluation.
Although our method operates directly on point clouds, point clouds cannot represent continuous regions in space, which makes direct evaluation using traditional IoU challenging. Therefore, we voxelized our results and used traditional IoU for scene completion and mIoU for semantic scene completion evaluation. While this introduces quantization error and potentially degrades our model’s performance, it aligns with the baseline settings and preserves their performance for a fair comparison.
IV-B Main Results
Based on the experimental setting described above, we present the results in Tab. I. Our results include the direct predictions from DiffSSC and the outcomes after refinement. We do the ablation study without the diffusion model, which means directly refining the output of Cylinder3D. We also report the results of baselines as a comparison.
@@ IV-B Main Results
TABLE I: Quantitative results on SemanticKITTI and SSCBench-KITTI360
Method Reference SemanticKITTI SSCBench-KITTI360
IoU(SC) mIoU(SSC) IoU(SC) mIoU(SSC)
LMSCNet [12] 3DV’20 48.24 15.43 33.64 13.47
JS3C-Net [15] AAAI’21 51.32 21.38 35.57 16.95
LODE [34] ICRA’23 50.61 18.22 38.24 15.39
Cylinder3D(refined) - 23.36 7.63 20.66 7.21
DiffSSC - 49.38 22.67 36.76 17.34
DiffSSC(refined) - 57.03 25.87 40.72 18.51
Best and second best results are highlighted.
@@
Figure 4: Qualitative results on SemanticKITTI and SSCBench-KITTI360. The 19 classes are shown without empty spaces. The estimated points that are located at the unknown region are visualized with $20\%$ opacity.
Qualitative results are presented in Fig.4. To emphasize the advantages of our approach, which operates directly on point clouds, we visualize samples from both the SemanticKITTI and SSCBench-KITTI360 datasets in point cloud form. For voxel-based methods, the point cloud is generated by sampling the center point of each occupied voxel. As shown in Fig.4, our DiffSSC model predicts more accurate semantic segmentation of the background and provides a more precise representation of the foreground shapes. Furthermore, the voxel-based baselines, which estimate the scene using two halves of a LiDAR sweep, exhibit discontinuous predictions at the boundary between the front and rear segments.
IV-C Model Analysis
(a) Noise schedule
(b) Ratio of global regularization
Figure 5: Hyperparameter study for the model’s performance
IV-C1 Noise Schedule
As mentioned in Sec.III, the noise schedule determines the intensity of noise added at each step, commonly including linear, cosine, and sigmoid schedules. We conducted a series of experiments to identify the most effective noise schedule for the SSC task. In Fig.5(a), we present the training curves sampled at each epoch, highlighting the convergence patterns for each schedule. Additionally, we compare the output of DiffSSC without refinement using each noise schedule in Tab.II, providing insights into their impact. The linear schedule, the simplest, was primarily used in early research. It shows slow and stable convergence, and its performance is significantly lower than that of the other two schedules. The cosine schedule, an improved function, introduces noise more gradually at the beginning and end, with a faster increase in the middle, balancing faster convergence with high final generation quality. The sigmoid schedule shares similarities with the cosine schedule, featuring an S-shaped curve but offering more precise control over the noise introduction, theoretically providing even greater potential. As a result, the cosine schedule converges significantly faster than the linear one. Although the sigmoid schedule does not converge as quickly as the cosine schedule, it is still noticeably faster than the linear schedule. They perform comparably, though the sigmoid schedule is slightly weaker than the cosine schedule. Therefore, in our main results, we adopted the cosine schedule.
TABLE II: Model’s performance based on different noise schedule functions
@@ IV-C1 Noise Schedule
Noise Schedule SemanticKITTI SSCBench-KITTI360
IoU(SC) mIoU(SSC) IoU(SC) mIoU(SSC)
Linear 45.26 19.02 33.32 14.99
Sigmoid 48.29 22.48 36.24 16.55
Cosine 49.38 22.67 36.76 17.34
Best and second best results are highlighted.
IV-C2 Regularization
We also investigated the model’s performance on SemanticKITTI under different ratios of global regularization, as shown in Fig. 5(b). When $\lambda=0$, indicating the use of only local $L_{2}$ regularization without global regularization, the model exhibited the worst performance, highlighting the benefits of incorporating global regularization. As the ratio increases, the model’s performance improves, peaking at $\lambda=5$ before declining. This suggests that excessive global regularization can constrain the model’s ability to generate finer details.

### v3

@@ III-C Denoiser Architecture
Figure 3: Architecture of our MinkUNet Denoiser. As shown in the red area, we design information fusion layers and insert them between MinkUNet blocks to integrate the conditional input and step information, guiding the generation of the point cloud.
@@
As shown in Fig. 3, the denoiser is based on the MinkUNet architecture [21, 22, 23, 24]. Given the feature $\mathcal{F}$ extracted from a layer of MinkUNet, we integrate the conditional input and step information between layers to obtain the fused feature $\mathcal{F}^{\prime}$. The raw semantic point cloud $\mathcal{X}$ is encoded as a conditional input $\mathcal{C}$. To embed the most relevant conditional input into the feature space, a closest point algorithm is employed to effectively align the conditional input with the features. Simultaneously, the step $t$ is encoded as $\tau$ using sinusoidal positional encodings. After passing through an MLP individually, the conditional input and step information are concatenated to form the weight $\mathcal{W}$. To align the dimensions with the feature $\mathcal{F}$, $\mathcal{W}$ is processed through an MLP to produce $\mathcal{W}^{\prime}$. Finally, $\mathcal{W}^{\prime}$ and $\mathcal{F}$ are element-wise multiplied to form the refined feature $\mathcal{F}^{\prime}$, which is then passed to the next layer.
@@ III-D Refinement
Inspired by Lyu et al. [28], we design a refinement and upsampling scheme based on MinkUNet to further enhance the density of the diffusion model’s output. This module predicts $k$ bias $b_{k}\in\mathbb{R}^{3}$ for each point position in the completed scene, while the semantics are propagated to the biased points. The refinement module offers a marginal improvement in scene quality, but it functions as interpolating points in the gaps, rather than learning to predict missing geometry and semantics. The main contribution is made by the diffusion model, as will be demonstrated in the ablation study.
IV Experiments
IV-A Benchmark Result
To conduct a comprehensive comparison with other leading SSC methods, we first evaluate our model on the SemanticKITTI [27] Benchmark. SemanticKITTI is a widely used autonomous driving dataset that provides point-wise annotations on raw LiDAR point clouds, extending the original KITTI dataset for semantic understanding tasks. The SSC Benchmark is a subtask within SemanticKITTI, focusing on predicting both the semantic class and occupancy status of each voxel within a grid volume. To obtain the ground truth, the annotated sequential scans are first accumulated and then chopped based on a predefined range represented in the LiDAR sensor’s coordinate system: $V_{\text{kitti}}=\{(x,y,z)\mid x\in[0,51.2]\text{ m},y\in[-25.6,+25.6]\text{ m},z\in[-3.2,+3.2]\text{ m}\}$. The extracted region is then voxelized into a 256×256×32256$\times$ 256$\times$ 32256 × 256 × 32 grid volume, where each voxel represents a $0.2^{3}\,\text{m}^{3}$ cube in the real world. Although our method operates directly on point clouds, the discrete nature of point clouds makes it challenging to directly evaluate performance using traditional IoU metrics, which are designed for continuous spatial regions. To address this, we voxelized our results and employed IoU for scene completion and mIoU for semantic scene completion evaluation. While voxelization introduces quantization errors that may slightly degrade our model’s performance, this approach ensures a fair and meaningful comparison.
TABLE I: Quantitative results on the SemanticKITTI benchmark.
[…]
LMSCNet [8] 3DV’20 55.3 17.0
S3CNet [12] CoRL’21 45.0 29.5
SSA-SC [29] IROS’21 58.8 23.5
JS3C-Net [30] AAAI’21 56.6 23.8
LODE [31] ICRA’23 51.2 23.4
SCPNet [32] CVPR’23 56.1 36.7
TALoS [33] NeurIPS’24 60.2 37.9
Ours - 63.4 27.4
IoU is used to evaluate only the occupancy status, while mIoU assesses performance across all semantic classes. Best and second best results are highlighted.
We follow the official dataset split: sequences 00-07 and 09-10 are used for training, 08 for validation, and 11-21 for testing. Model performance is evaluated through the official online benchmark server. The model is trained on an NVIDIA A6000 GPU for 20 epochs. For the diffusion parameters, we employ a cosine schedule to modulate the intensity of noise at each step. Specifically, we set $\beta_{0}=3.5\times 10^{-5}$ and $\beta_{T}=0.007$, with the number of diffusion steps $T=1000$, and define $\beta_{1},\dots,\beta_{T-1}$ using the following equation.
@@ IV-A Benchmark Result
We set the ratio of global regularization to $\lambda_{p}=5.0$ and $\lambda_{s}=4.0$. Additionally, we define the scaling factors as $\sigma_{p}=1.0$ and $\sigma_{s}=0.2$. As shown in Tab. I, our method outperforms all state-of-the-art LiDAR-based approaches in scene completion, indicating that the model effectively captures geometric information. However, for semantic scene completion (measured by mIoU), a obvious gap remains compared to SCPNet and TALoS. It is important to note that both methods leverage additional information beyond single-frame input. Specifically, SCPNet utilizes multi-frame knowledge distillation to enhance the prediction of the current frame, while TALoS employs a test-time adaptation strategy, incorporating multi-frame observations and future data to refine its outputs. In contrast, DiffSSC relies solely on single-frame information and does not incorporate online model adjustments.
While the SemanticKITTI SSC benchmark provides a valuable framework for evaluating SSC models, its design is notably influenced by early indoor SSC tasks. These early methods mainly relied on RGB-D cameras, which inherently limited perception to front-facing scenes. To leverage the technical foundations of these methods and simplify their transition to outdoor environments, SemanticKITTI restricts the LiDAR field of view to the front half, covering only the range $[-90^{\circ},+90^{\circ}]$ in the sensor coordinate system, while completely ignoring the rear part of the scene. Although this simplification facilitated early exploration of outdoor SSC, it fundamentally deviates from the natural properties of LiDAR data and the core requirements of autonomous driving tasks. Rear-view perception is equally critical for driving safety, especially for tasks like lane changes, reversing, and obstacle avoidance. Since LiDAR sensors capture data through continuous $360^{\circ}$ rotations without directional bias, restricting input to the front half disrupts the spatial consistency of the data and limits the model’s ability to fully understand the driving environment. To address this limitation, we extend the scene completion task to a full $360^{\circ}$ panoramic view, preserving the intrinsic characteristics of LiDAR data and providing a more comprehensive representation of real-world autonomous driving scenarios.
IV-B Extended Experiment on Panoramic Scenarios
IV-B1 Panoramic Settings
In the raw SSC setting of SemanticKITTI, the scene is limited to a cuboid region $V_{\text{kitti}}$, covering the range $[-90^{\circ},+90^{\circ}]$. To extend this to the full $[-180^{\circ},+180^{\circ}]$ panoramic range while preserving spatial symmetry, the panoramic volume is defined as the combination of two SemanticKITTI volumes facing forward and backward. Specifically, in the LiDAR’s local coordinate system, this is formulated as: $V_{\text{pano}}=\{(x,y,z)\mid x\in[-51.2,51.2]\text{ m},y\in[-25.6,+25.6]\text{ m},z\in[-3.2,+3.2]\text{ m}\}$. The front and rear halves of the panoramic volume are designed to be identical to the original SemanticKITTI volume. Given the similar statistical characteristics of LiDAR data in the front and rear regions, retaining the original SemanticKITTI volume settings allows baseline methods to be seamlessly transferred to the panoramic setting without significant performance degradation.
We generate the ground truth following the guidelines of SemanticKITTI. First, using the pose information of each frame, we construct a global map by aggregating the semantic LiDAR sweeps within the sequence. Next, we extract the region within $V_{\text{pano}}$, with the LiDAR positioned at its center. Additionally, unknown areas defined by the raw dataset are mapped into $V_{\text{pano}}$ using the known poses, and these regions are excluded from the evaluation.
Besides SemanticKITTI, we also conduct panoramic experiments on the SSCBench-KITTI360 dataset [34]. SSCBench-KITTI360 is another SSC benchmark derived from KITTI-360 [35], with semantic information aligned to the SemanticKITTI format. This consistency allows SSC methods evaluated on SemanticKITTI to be seamlessly transferred to the KITTI-360 scenario. However, both benchmarks only use the front half of the LiDAR scan as input. To address this, we apply the same panoramic processing approach to SSCBench-KITTI360 as we did for SemanticKITTI.
IV-B2 Baselines
@@
We compare our approach against LMSCNet [8], JS3C-Net [30], and LODE [31]. Both LMSCNet and JS3C-Net take the front half of the quantized LiDAR sweep as input and are evaluated on the SSC benchmark of SemanticKITTI. LODE primarily focuses on geometry completion using implicit representations; however, to demonstrate its flexibility, the authors also report results with extended semantic parsing.
To enable these baselines to predict panoramic scenes, we split the $360^{\circ}$ LiDAR point cloud into two halves and feed them separately into the baseline models. The front half follows the same settings as in SemanticKITTI, while the rear half is rotated by $180^{\circ}$ before being passed to the models. After obtaining predictions for both halves, we concatenate them to form the complete panoramic scene. Although the baselines were trained solely on the front part of the scene, the statistical characteristics of LiDAR data in the front and rear regions are similar. This suggests that models trained on the front half remain effective when applied to the rear region. This approach minimizes performance degradation due to domain shift, ensuring a fair comparison. Additionally, we directly utilized the official code and pretrained checkpoints from the baselines to predict the panoramic scenes, further maintaining consistency in evaluation.
While these baselines have reported results on SemanticKITTI, they had not previously been tested on SSCBench-KITTI360 [34]. To supplement our evaluation, we ran these baselines on SSCBench-KITTI360 without fine-tuning. Since the semantic labels and the overall pipeline in SSCBench-KITTI360 are consistent with SemanticKITTI, the baselines could be seamlessly applied to this dataset.
IV-B3 Training and Inference
Since our method operates directly on point clouds rather than voxel-based volumes, the processing of training pairs differs from that of the baselines. Although we generate the ground truth by aggregating the semantic LiDAR sweeps and extracting point clouds within the $V_{\text{pano}}$, we do not further voxelize the data. To facilitate better diffusion and learning, we define the generated scene range during both training and inference as a spherical area centered on the LiDAR with a radius of 60 meters, i.e., $V_{\text{sphere}}=\{(x,y,z)\mid\sqrt{x^{2}+y^{2}+z^{2}}\leq 60\,\text{m}\}$. During evaluation, we voxelize the predicted point cloud scene and extract the portion within $V_{\text{pano}}$. Thus, while the generated scene range differs from $V_{\text{pano}}$, the evaluation region is restricted to $V_{\text{pano}}$, ensuring consistency with the baselines. Moreover, the model does not leverage any information outside $V_{\text{pano}}$ but within $V_{\text{sphere}}$ during the learning process, ensuring a fair comparison with the baselines. Our model is trained and validated purely on SemanticKITTI, using sequences 00-07 for training and sequences 09-10 for validation. We evaluate our model on the official validation sets of both datasets: sequence 08 of SemanticKITTI and sequence 07 of SSCBench-KITTI360. Since the baselines were not fine-tuned on SSCBench-KITTI360, we likewise did not perform any fine-tuning on SSCBench-KITTI360.
IV-B4 Experimental Results
Based on the experimental settings described above, we compare the performance of existing SSC methods with our approach in Tab. II. Although voxel-based evaluation introduces quantization errors, the output of our diffusion model surpasses all baselines, demonstrating the effectiveness of our method.
@@ IV-B4 Experimental Results
TABLE II: Quantitative results of the panoramic experiments on the SemanticKITTI and SSCBench-KITTI360 validation sets.
Method SemanticKITTI SSCBench-KITTI360
[…]
LMSCNet [8] 48.2 15.4 33.6 13.5
JS3C-Net [30] 51.3 21.4 35.6 17.0
LODE [31] 50.6 18.2 38.2 15.4
Ours 60.3 26.7 47.3 20.4
Best results are highlighted in bold.
Qualitative results are presented in Fig. 4. To highlight the advantages of our approach, which operates directly on point clouds, we visualize samples from both the SemanticKITTI and SSCBench-KITTI360 datasets in point cloud form. For voxel-based methods, point clouds are generated by sampling the center point of each occupied voxel. As shown in Fig. 4, our DiffSSC model predicts more accurate semantic segmentation of the background and offers a more precise representation of foreground shapes. Moreover, voxel-based baselines, which estimate the scene using two halves of a LiDAR sweep, exhibit discontinuities at the boundary between the front and rear parts. This further underscores the importance of learning from panoramic LiDAR data.
@@
[…]
IV-C Ablation Studies
TABLE III: The results of ablation studies for the proposed DiffSSC. The performance is evaluated on SemanticKITTI and SSCBench-KITTI360 validation sets. Best results are highlighted.
@@ IV-C Ablation Studies
Method Module Noise Schedule Regularization SemanticKITTI SSCBench-KITTI360
[…]
Ours A $\checkmark$ $\checkmark$ $\checkmark$ - - $\checkmark$ $\checkmark$ $\checkmark$ 60.3 26.7 47.3 20.4
Module-level B $\checkmark$ - $\checkmark$ - - - - - 23.4 7.6 20.7 7.2
C $\checkmark$ $\checkmark$ - - - $\checkmark$ $\checkmark$ $\checkmark$ 58.7 26.3 42.1 19.3
Policy-level D $\checkmark$ $\checkmark$ - $\checkmark$ - - $\checkmark$ $\checkmark$ 52.6 20.7 37.1 16.3
E $\checkmark$ $\checkmark$ - - $\checkmark$ - $\checkmark$ $\checkmark$ 56.9 25.8 40.9 18.5
F $\checkmark$ $\checkmark$ - - - $\checkmark$ $\checkmark$ - 40.3 10.9 30.7 9.6
To systematically analyze the contribution of each component in our model, particularly the core diffusion model and key learning strategies, we ablate our method on SemanticKITTI and SSCBench-KITTI360, with results summarized in Tab. III. Our model architecture comprises three primary modules: semantic segmentation, diffusion, and refinement. We performed module-level ablations to evaluate the individual contributions of each submodule. Additionally, we conducted policy-level ablations to examine the influence of two critical factors on the diffusion model: the noise schedule function and global regularization.
IV-C1 Module-level
The semantic segmentation module is essential for providing initial semantic priors. Therefore, we focused on analyzing the individual contributions of the diffusion and refinement modules. In Method B, we removed the core diffusion module from our pipeline, directly refining the output of Cylinder3D. Without the diffusion step, the noise schedule and regularization become irrelevant. This resulted in poor performance, indicating that refinement alone cannot effectively predict unknown areas in the scene. In Method C, we used only the diffusion model without refinement. While performance slightly dropped compared to the full Method A, this suggests that refinement improves scene densification but relies on accurate predictions from the diffusion model.
IV-C2 Policy-level
As mentioned in Sec. III, the noise schedule determines the intensity of noise injected at each step, commonly including linear, cosine, and sigmoid schedules. We investigated the impact of different noise schedules on the diffusion process. To isolate this effect, we removed the refinement module, allowing a clearer view of how the noise schedule influences diffusion. Results are shown in Tab. III. Comparing Method C, D, and E, we observe that the linear schedule, being the simplest form, performs significantly worse than the other two schedules. The cosine schedule, with its S-shaped curve and precise control over noise introduction, balances faster convergence with high final generation quality, achieving the best results. The sigmoid schedule also shows competitive performance, slightly lagging behind the cosine schedule. Therefore, we adopted the cosine schedule in our main results.
We also investigated the impact of global regularization on model performance. By setting $\lambda_{p}=5.0$ and $\lambda_{s}=4.0$, we removed global regularization in Method F. Compared to Method C ($\lambda_{p}=\lambda_{s}=0$), the model exhibited poorer performance, highlighting the advantages of incorporating global regularization.